OpenStudy (anonymous):

Find the particular solution of the differential equation (x^2/y^2-3)(dy/dx)=1/2y satisfying the initial condition y(1)=sqrt(4). solve for y. your answer should be a function of x.

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{1}{x^{2}} dx = 2 \int\limits_{}^{} \frac{1}{y(y^2-3)} dy\] are you familiar with partial fractions decomposition?

myininaya (myininaya):

it looks like separation of variables can be done

OpenStudy (anonymous):

not really. i know the formula for exponential decay

OpenStudy (anonymous):

in order to separate variables, we need to have all y's or all x's right

myininaya (myininaya):

no we need to be able to write f(y) dy =g(x) dx to use separation of variables

myininaya (myininaya):

john already set it up gj john

OpenStudy (anonymous):

ok. where do i go from there?

myininaya (myininaya):

use partial fractions for right hand side

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{2}{y(y^{2}-3}} = \frac{A}{y} + \frac{B}{y-\sqrt{3}} +\frac{C}{y+\sqrt{3}\]

OpenStudy (anonymous):

bah

OpenStudy (anonymous):

use A/y + B/y-sqrt(3) + C/y+sqrt(3)

OpenStudy (anonymous):

2= A (y^2-3) + B y (y+sqrt3) + C y (y-sqrt3) and find A,B and C by plugging in y=0 sqrt(3) and - sqrt(3)

OpenStudy (phi):

isn't the 2y in the numerator??

OpenStudy (phi):

assuming 1/2y is really 1/(2y)

OpenStudy (anonymous):

if it is, that makes life so much easier :)

OpenStudy (phi):

Yes

myininaya (myininaya):

lol

OpenStudy (anonymous):

\[-\frac{1}{x}+C = \ln(y^2-3) \]

OpenStudy (anonymous):

solve for C?

OpenStudy (anonymous):

\[y= \sqrt{e^{-1/x+c} +3}\]

OpenStudy (phi):

@bbaker plug in x=1 and y=2 and solve for C

OpenStudy (anonymous):

4=e^-1 * e^c+3...............0=-1+C and C=1

myininaya (myininaya):

use the equation before the second one john wrote i think it is much easier to see what C is in that equation if x=1 and y=2

myininaya (myininaya):

-1/x+c=ln(y^2-3) -1/1+c=ln(2^2-3) -1+c=ln(1) -1+c=0 c=1

OpenStudy (anonymous):

if c=1 then y= 3+e^-1/x+1?

myininaya (myininaya):

if c=1, we can write -1/x+1=ln(y^2-3)

myininaya (myininaya):

we should put absolute value around that y^2-3 by the way

OpenStudy (phi):

Though typically people use the exponential form. I think you should add some parens to make your answer more clear. and put a square root around it

OpenStudy (anonymous):

from there solve for y?

myininaya (myininaya):

e^(-1/x+1)=y^2-3 3+e^(-1/x+1)=y^2 take square root of both sides don't forget to write plus or minus

OpenStudy (anonymous):

so as a function of x y=sqrt(3+e^(-1/x+1)) or -sqrt(3+e^(-1/x+1))